# Investigating Large Language Models for Code Vulnerability Detection: An Experimental Study

### v1

@@ Abstract
keywords:
Code Vulnerability Detection, Large Language Model, Code Intelligence, Cyber Security, Experimental Study
1 Introduction
@@
The graph-based models, represented by Devign [2], attempt to transform the source code into the code graph, extract the code elements as graph nodes, and analyze vulnerabilities through graph representation learning. The sequence-based models, represented by CodeBERT [3], aim to regard the source code as a sequence of tokens, and utilize RNN-based or more advanced Transformer-based pre-trained language models to capture the vulnerable pattern within the code. The graph-based models are good at capturing the structural information of the code but struggle to capture long-distance association among the nodes, especially when the code content gets larger. Meanwhile, recent studies [4, 5] and our fine-grained statistics across five commonly-used datasets in Table 2 point out that the vulnerable code pattern tends to exist in the long code context. Thus, more efforts are put to the sequence-based models to detect code vulnerabilities, especially the pre-trained language models [3].
@@
In this work, to bridge the above existing gap, we provide an early experimental investigation on fine-tuning LLMs on the CVD datasets, particularly focusing on 4 widely-used open-source Llama models, including two rarely evaluated LLMs (i.e. Llama-3 and Llama-3.1 [35]). We revisit related literature, choose 5 most commonly-used CVD datasets, and additionally integrate 3 graph-based models and 2 medium-size BERT based sequence models into a unified codebase. We also study the impacts of class imbalance and code sequence length to the model performance with quantitative experiments. In addition, all source codes are open-source to facilitate related communities.
@@
  • We conduct a systematic investigation into the capabilities of fine-tuned LLMs for code vulnerability detection. Through comprehensive experiments and analysis, we evaluate the performance of 4 LLMs across 5 distinct code vulnerability datasets, involving the largest number of datasets among existing empirical studies. The number of involved datasets is the largest among existing empirical studies. Furthermore, we compare their effectiveness with 3 representative graph-based model and 2 medium-size pre-trained sequence models.
@@
The remainder of this paper is organized as below. Section 2 discusses the CVD task and three kinds of model architecture to tackle this task. Section 3 states our motivation to carry out this work. Section 4 introduces the problem definition and related preliminary knowledge. Section 5 elaborates on our evaluated models and pre-processed experimental datasets. Section 6 showcases the experimental results and related findings. Section 7 summarizes this study, then discusses the limitation of this work and potential future directions.
2 Related Works
@@
Medium-size sequence models. Some early attempts like VulDeePecker [52] and SeSyVR [51] aim to identify code vulnerability via light-weight sequence models like TextCNN, RNN or LSTM [38, 13]. However, these early small-size sequence models are quickly surpassed by pre-trained Transformer [30] based language models which are widely trained on a large corpus [20]. The parallel processing ability of these traditional sequence models is also limited, and it is also difficult to capture the association between long-distance tokens. Meanwhile, Transformer-based pre-trained language models [30, 31] (or code pre-trained models as referred in [17]) have better scalability to the input context length than these early sequence models. Achieving such input length scalability and long-distance token association is mainly credited to the attention mechanism [30]. These pre-trained language models often adopt a new learning paradigm of ’pre-training and fine-tuning’, where the pre-training stage aims to learn general semantic information on large-scale corpus and the fine-tuning stage aims to relatively quickly adapt to the downstream CVD tasks [6]. Representative models include CodeBERT [3] and UniXcoder [48], which will be detailedly discussed in Section 5.3
@@ 3 Motivation
Unified and Easy-to-use Open-source Implementation. In addition, during we carry out this study, we find there lacks a unified open-source codebase to train and evaluate both graph-based models, medium-size sequence models and LLMs, which brings obstacles for related communities to carry out re-implementation. Therefore, based on their open-source Github or HuggingFace repositories listed in Section 5.1, we implement nine related models as shown in Table 1 and carefully integrate them into one unified codebase to better facilitate related communities. For LLMs, we investigate two advanced Llama series (Llama-3 and Llama-3.1 [35]) which are rarely studied in previous CVD works. Meanwhile, we provide the five most commonly used pre-processed datasets with a unified format. We organize all training or fine-tuning codes in an easy-to-use manner, which make it easier for re-implementations of graph-based models, medium-size sequence models and LLMs.
4 Preliminaries
4.1 Problem Definition
@@ 4.3 Instruction Tuning
5 Benchmark Design
5.1 Evaluated Models
@@
Table 1: Details on the evaluated models.
Model Arch. Venue Scale Type Main Model Component(s)
@@ 5.2 CVD Datasets
We summarize related statistics of these datasets in Table 2. Except Devign [2], other datasets exhibit obvious class imbalance. We subsample part of the samples in Draper, BigVul and diverseVul. More details regarding our pre-processing procedures can be found in Section 5.3.
@@
Table 2: Details on the evaluated CVD datasets. Vul. Ratio indicates the proportion of the vulnerable samples across the samples ($\sim$50indicates a relatively balanced dataset). The number of samples before subsampling is indicated in parenthesis; see Section 5.3 for details.
@@ 5.3 Dataset Pre-processing
As fine-grained statistics listed in Section 5.2, our study involves 5 influential and widely-used datasets in this field. This enables us to observe how well each model performs when confronted with various types of vulnerabilities. Due to the significant number of involved datasets and the obvious differences in attributes such as sample size and positive sample ratios, we applied the following data pre-processing steps in the main experiments to ensure fairness in evaluation:
  • Filtering. As mentioned in Section 5.2, the quality of code datasets varies significantly. We find anomalies in some samples during data preprocessing. In the DiverseVul dataset, We are unable to trace some samples based on their ‘project’ and ‘commit_id’ attributes. In the Draper dataset, annotation inaccuracies are prevalent, particularly in code samples associated with multiple CWE types, where obvious labeling errors are found. To address this, we perform an initial filtering of these two datasets to exclude anomalous samples. Note that the data quality and label noise issues are also pointed out in previous works [22], which leaves space for future works.
  • Formatting. Diverse representations and storage formats of samples pose challenges for conducting unified experiments. We format every dataset in order to avoid this. We assign a unique index to each sample and used the ‘code’ and ‘label’ attributes to represent every sample’s code and label respectively. Additionally, we retain additional attributes specific to each dataset, such as CWE type and commit ID, which can assist with future works like vulnerability line extraction and vulnerability classification. Notably, only the ‘code’ and ‘label’ attributes are used in all of our experiments in Section 6. To facilitate the fine-tuning of LLMs to adapt the CVD classification task, the labels are annotated to $\mathbf{1}$ or $\mathbf{0}$ to denote the vulnerable and benign class, following the previous successful practice [16, 17]. Data are formatted using the general instruction fine-tuning template provided by Alpaca [58] format, as illustrated in Figure 1. In this template, [Input] and [Output] are derived from the aforementioned data preparation process, while [Task Prompt] guides the LLM to generate task-specific outputs based on different tasks.
@@
6 Experiments
6.1 Experimental Settings
@@
Datasets We conduct experiments on 5 widely-used code vulnerability datasets as elaborated in Section 5.2. We divided each dataset into two subsets (long samples and short samples) based on the sample length. Subsequently, each subset is split into train, validation, and test sets in the ratio of 8:1:1. Graph models cannot directly process sequential data. Therefore, we used Joern [65], a CPG [61] based C/C++ code analysis tool, to convert each sequence sample into a code graph, which is aligned with the processing methods of Devign [2] and ReVeal [24]. To obtain feature vectors for each node in the graph, we trained a Word2Vec [64] model with a vector size of 200 to vectorize each token. The converted graph dataset is stored in JSON format.
Models and Hyper-parameters We evaluate 3 graph-based models, 2 medium-size sequence models and 4 large language models (LLMs). We provide the Github repositories for all implementations of fine-tuning LLMs and baseline models in Section 5.1. For the classical graph-based model Devign[2], the input feature size and graph embedding size are set to 200. Adam is used as the optimizer with a learning rate of 1e-4 and a weight decay of 1e-3. Both of medium-size sequence models and all graph-based models except Devign are methods based on Transformer encoder, and their hyperparameter settings are consistent. The block size of them is set to 512 for short sample datasets and 1024 for long sample datasets. AdamW is used as the optimizer with a learning rate of 2e-5. For fine-tuning LLMs, we configure the model parameters using the default settings provided by Meta AI®. We employ LoRA[21] for fine-tuning, setting the rank to 16, the scaling factor $\alpha$ to 32, and dropout rate to 0.05. AdamW is used as the optimizer with a learning rate of 1e-4, and the model is trained for 5 epoch. We fine-tune the $q_{proj}$, $v_{proj}$, $k_{proj}$, and $o_{proj}$ weight matrices in the self-attention layers following [17].
@@
   In highly imbalanced CVD datasets as introduced in Section 5.2, the commonly used accuracy (Acc.) metric yields misleadingly high performances that result from systematically predicting the majority class [9]. Therefore, F1-Score can be more precise for the CVD task which is our main technical metric. In addition, FPR can assist to understand why a given classification system underperforms.
Table 3: Main experimental metrics (%). We use the F1-Score as the main analyzed metric. The bold denotes the best result on this dataset while the underlined denotes the second place result on this dataset. - denotes we do not conduct related experiments because these models do not support samples with more than 512 tokens as discussed in Section 5.1.
@@
6.2 Analysis on Main Experiments
For main experiments, we train and evaluate 3 graph-based models, 2 medium-size sequence models and 4 LLMs on 5 widely-used code vulnerability datasets, as introduced in Section 5. Notably, only Devign is a relatively class balanced dataset, and other four datasets exhibit obvious class imbalance as shown in Table 2.
GraphCodeBERT, CodeBERT and UniXcoder cannot be evaluated on long sample datasets because they are all based on the RoBERTa [68] architecture. The learnable position encoding layer of RoBERTa limits the input sequence length to 512 [6], meaning that long samples will be truncated to 512. Although ReGVD is also based on the RoBERTa architecture, it only uses the pretrained embedding layer and does not involve position encoding or any subsequent layers. Therefore, we can evaluate ReGVD on long sample datasets. Table 3 provides a detailed summary of the main experimental results.
Finding 1: The performance of all LLMs and other models tend to be influenced by the class imbalance of the dataset. All models perform significantly better on the Devign dataset than on the other datasets, and Devign is also the most balanced dataset with nearly equal numbers of positive and negative samples. In contrast, all models tend to perform poorly on the DiverseVul dataset, which has the fewest positive samples. The difference between these two datasets is most clearly reflected in recall. It is worth noting that in some experiments, all metrics except accuracy are 0. This phenomenon is most commonly observed on ReGVD, which only present normal metrics on the most balanced Devign dataset. Moreover, both Llama-2 and CodeLlama showed this anomaly on the most imbalanced DiverseVul dataset. It emphasizes the important role of data balance.
Finding 2: The medium-size sequence models excel on short sample datasets, generally outperforming LLMs. On the ReVeal, Draper, and DiverseVul datasets, both of the medium-size sequence models achieve the highest and second-highest F1-scores respectively. There is an evident performance difference between the LLM and medium-sized sequence models. Although they have generally similar precisions, LLMs’ recalls are significantly lower than the medium-sized sequence models, resulting in lower F1- scores. This gap narrows as the dataset becomes more balanced, with the Devign dataset showing the smallest gap. We conclude that medium-sized sequence models are less affected by a low proportion of positive samples than the LLMs. The code pre-trained Transformer encoder enables them to capture vulnerability features more accurately even with limited vulnerability data.
@@ 6.2 Analysis on Main Experiments
From the above analysis, it is clear that the proportion of positive samples in the dataset plays a decisive role in the CVD performance of trained models, while the impact of sample length should not be overlooked. To specifically investigate the effects of positive sample ratio and length on fine-tuning LLMs, we have designed experiments in Section 6.3 and 6.4 respectively.
6.3 Analysis on Datasets with Varying Postive Sample Ratios
As two recent study [4, 5] and our statistics in Table 2 point out, long-tailed distribution within CVD datasets could pose a challenge for LLMs-based vulnerability detection solutions, and we can also observe this in our main experiments in Table 3. Thus, we carry out the re-sampling experiments which creates more balanced datasets. We subsample the Draper and DiverseVul datasets [26, 22] to make the positive samples (i.e. vulnerable) to occupy more percentage across the training dataset, and the ratio is incrementally set to 10%, 20%, 30%, 40% and 50%. Similar to the main experiment, the size of each sampled dataset is controlled at 25,000. We select CodeBERT as the studied medium-size sequence model and the LLama-3.1 as the studied LLM.
@@ 6.3 Analysis on Datasets with Varying Postive Sample Ratios
6.4 Analysis on Analysis on Datasets with Varying Sample Lengths

### v2

@@ Abstract
Index Terms:
Code Vulnerability Detection, Large Language Model, Code Intelligence, Cyber Security, Experimental Study.
I Introduction
@@
The graph-based models, represented by Devign [2], attempt to transform the source code into the code graph, extract the code elements as graph nodes, and analyze vulnerabilities through graph representation learning. The sequence-based models, represented by CodeBERT [3], aim to regard the source code as a sequence of tokens, and utilize RNN-based or more advanced Transformer-based pre-trained language models to capture the vulnerable pattern within the code. The graph-based models are good at capturing the structural information of the code but struggle to capture long-distance association among the nodes, especially when the code content gets larger. Meanwhile, recent studies [4, 5] and our fine-grained statistics across five commonly-used datasets in Table II point out that the vulnerable code pattern tends to exist in the long code context. Thus, more efforts are put to the sequence-based models to detect code vulnerabilities, especially the pre-trained language models [3].
@@
In this work, to bridge the above existing gap, we provide an early experimental investigation on fine-tuning LLMs on the CVD datasets, particularly focusing on 4 widely-used open-source Llama-series models, including two rarely evaluated LLMs (i.e. Llama-3 and Llama-3.1 [35]). We revisit related literature, choose 5 most commonly-used CVD datasets, and additionally integrate 3 graph-based models and 2 medium-size BERT based sequence models into a unified codebase. We also study the impacts of class imbalance and code sequence length to the model performance with quantitative experiments. In addition, all source codes with clear hand-on guidance are already open-source to facilitate related communities for more convenient reproduction of corresponding models.
@@
  • We conduct a systematic investigation into the capabilities of fine-tuned LLMs for code vulnerability detection. Through comprehensive experiments and analysis, we evaluate the performance of 4 LLMs across 5 distinct code vulnerability datasets, involving the largest number of datasets among existing empirical studies. Furthermore, we compare their effectiveness with 3 representative graph-based model and 2 medium-size pre-trained sequence models.
@@
The remainder of this paper is organized as below. Section II discusses the CVD task and three kinds of model architecture to tackle this task. Section III states our motivation to carry out this work. Section IV introduces the problem definition and related preliminary knowledge. Section V elaborates on our evaluated models and pre-processed experimental datasets. Section VI showcases the experimental results and related findings. Section VII summarizes this study, then discusses the limitation of this work and potential future directions.
II Related Works
@@
Medium-size sequence models. Some early attempts like VulDeePecker [53] and SeSyVR [52] aim to identify code vulnerability via light-weight sequence models like TextCNN, RNN or LSTM [38, 13]. However, these early small-size sequence models are quickly surpassed by pre-trained Transformer [30] based language models which are widely trained on a large corpus [20]. The parallel processing ability of these traditional sequence models is also limited, and it is also difficult to capture the association between long-distance tokens. Meanwhile, Transformer-based pre-trained language models [30, 31] (or code pre-trained models as referred in [17]) have better scalability to the input context length than these early sequence models. Achieving such input length scalability and long-distance token association is mainly credited to the attention mechanism [30]. These pre-trained language models often adopt a new learning paradigm of ’pre-training and fine-tuning’, where the pre-training stage aims to learn general semantic information on large-scale corpus and the fine-tuning stage aims to relatively quickly adapt to the downstream CVD tasks [6]. Representative models include CodeBERT [3] and UniXcoder [49], which will be detailedly discussed in Section V-C
@@ III Motivation
Unified and Easy-to-use Open-source Implementation. In addition, during we carry out this study, we find there lacks a unified open-source codebase to train and evaluate both graph-based models, medium-size sequence models and LLMs, which brings obstacles for related communities to carry out re-implementation. Therefore, based on their open-source Github or HuggingFace repositories listed in Section V-A, we implement nine related models as shown in Table I and carefully integrate them into one unified codebase to better facilitate related communities. For LLMs, we investigate two advanced Llama series (Llama-3 and Llama-3.1 [35]) which are rarely studied in previous CVD works. Meanwhile, we provide the five most commonly used pre-processed datasets with a unified format. We organize all training or fine-tuning codes in an easy-to-use manner, which make it easier for re-implementations of graph-based models, medium-size sequence models and LLMs.
IV Preliminaries
IV-A Problem Definition
@@ IV-C Instruction Tuning
V Benchmark Design
V-A Evaluated Models
@@
TABLE I: Details on the evaluated models.
Model Arch. Venue Parameter Scale Type Main Model Component(s)
@@ V-B CVD Datasets
We summarize related statistics of these datasets in Table II. Except Devign [2], other datasets exhibit obvious class imbalance. We subsample part of the samples in Draper, BigVul and diverseVul. More details regarding our pre-processing procedures can be found in Section V-C.
@@
TABLE II: Details on the evaluated CVD datasets. Vul. Ratio indicates the proportion of the vulnerable samples across the samples ($\sim$50indicates a relatively balanced dataset). The number of samples before subsampling is indicated in parenthesis; see Section V-C for details.
@@ V-C Dataset Pre-processing
As fine-grained statistics listed in Section V-B, our study involves 5 influential and widely-used datasets in this field. This enables us to observe how well each model performs when confronted with various types of vulnerabilities. Due to the significant number of involved datasets and the obvious differences in attributes such as sample size and positive sample ratios, we applied the following data pre-processing steps in the main experiments to ensure fairness in evaluation:
  • Filtering. As mentioned in Section V-B, the quality of code datasets varies significantly. We find anomalies in some samples during data preprocessing. In the DiverseVul dataset, We are unable to trace some samples based on their ‘project’ and ‘commit_id’ attributes. In the Draper dataset, annotation inaccuracies are prevalent, particularly in code samples associated with multiple CWE types, where obvious labeling errors are found. To address this, we perform an initial filtering of these two datasets to exclude anomalous samples. Note that the data quality and label noise issues are also pointed out in previous works [22], which leaves space for future works.
  • Formatting. Diverse representations and storage formats of samples pose challenges for conducting unified experiments. We format every dataset in order to avoid this. We assign a unique index to each sample and used the ‘code’ and ‘label’ attributes to represent every sample’s code and label respectively. Additionally, we retain additional attributes specific to each dataset, such as CWE type and commit ID, which can assist with future works like vulnerability line extraction and vulnerability classification. Notably, only the ‘code’ and ‘label’ attributes are used in all of our experiments in Section VI. To facilitate the fine-tuning of LLMs to adapt the CVD classification task, the labels are annotated to $\mathbf{1}$ or $\mathbf{0}$ to denote the vulnerable and benign class, following the previous successful practice [16, 17]. Data are formatted using the general instruction fine-tuning template provided by Alpaca [59] format, as illustrated in Figure 1. In this template, [Input] and [Output] are derived from the aforementioned data preparation process, while [Task Prompt] guides the LLM to generate task-specific outputs based on different tasks.
@@
VI Experiments
VI-A Experimental Settings
@@
Datasets We conduct experiments on 5 widely-used code vulnerability datasets as elaborated in Section V-B. We divided each dataset into two subsets (long samples and short samples) based on the sample length. Subsequently, each subset is split into train, validation, and test sets in the ratio of 8:1:1. Graph models cannot directly process sequential data. Therefore, we used Joern [66], a CPG [62] based C/C++ code analysis tool, to convert each sequence sample into a code graph, which is aligned with the processing methods of Devign [2] and ReVeal [24]. To obtain feature vectors for each node in the graph, we trained a Word2Vec [65] model with a vector size of 200 to vectorize each token. The converted graph dataset is stored in JSON format.
Models and Hyper-parameters We evaluate 3 graph-based models, 2 medium-size sequence models and 4 large language models (LLMs). We provide the Github repositories for all implementations of fine-tuning LLMs and baseline models in Section V-A. For the classical graph-based model Devign[2], the input feature size and graph embedding size are set to 200. Adam is used as the optimizer with a learning rate of 1e-4 and a weight decay of 1e-3. Both of medium-size sequence models and all graph-based models except Devign are methods based on Transformer encoder, and their hyperparameter settings are consistent. The block size of them is set to 512 for short sample datasets and 1024 for long sample datasets. AdamW is used as the optimizer with a learning rate of 2e-5. For fine-tuning LLMs, we configure the model parameters using the default settings provided by Meta AI®. We employ LoRA[21] for fine-tuning, setting the rank to 16, the scaling factor $\alpha$ to 32, and dropout rate to 0.05. AdamW is used as the optimizer with a learning rate of 1e-4, and the model is trained for 5 epoch. We fine-tune the $q_{proj}$, $v_{proj}$, $k_{proj}$, and $o_{proj}$ weight matrices in the self-attention layers following [17].
@@
   In highly imbalanced CVD datasets as introduced in Section V-B, the commonly used accuracy (Acc.) metric yields misleadingly high performances that result from systematically predicting the majority class [9]. Therefore, F1-Score can be more precise for the CVD task which is our main technical metric. In addition, FPR can assist to understand why a given classification system underperforms.
TABLE III: Main experimental metrics (%). We use the F1-Score as the main analyzed metric. The bold denotes the best result on this dataset while the underlined denotes the second place result on this dataset. - denotes we do not conduct related experiments because these models do not support samples with more than 512 tokens as discussed in Section V-A.
@@
VI-B Analysis on Main Experiments
For main experiments, we train and evaluate 3 graph-based models, 2 medium-size sequence models and 4 LLMs on 5 widely-used code vulnerability datasets, as introduced in Section V. Notably, only Devign is a relatively class balanced dataset, and other four datasets exhibit obvious class imbalance as shown in Table II.
GraphCodeBERT, CodeBERT and UniXcoder cannot be evaluated on long sample datasets because they are all based on the RoBERTa [69] architecture. The learnable position encoding layer of RoBERTa limits the input sequence length to 512 [6], meaning that long samples will be truncated to 512. Although ReGVD is also based on the RoBERTa architecture, it only uses the pretrained embedding layer and does not involve position encoding or any subsequent layers. Therefore, we can evaluate ReGVD on long sample datasets. Table III provides a detailed summary of the main experimental results.
Finding 1: The performance of all LLMs and other models tend to be influenced by the class imbalance of the dataset. All models perform significantly better on the Devign dataset than on the other datasets, and Devign is also the most balanced dataset with nearly equal numbers of positive and negative samples. In contrast, all models tend to perform poorly on the DiverseVul dataset, which has the fewest positive samples. The difference between these two datasets is most clearly reflected in recall. It is worth noting that in some experiments, all metrics except accuracy are 0. This phenomenon is most commonly observed on ReGVD, which only present normal metrics on the most balanced Devign dataset. Moreover, both Llama-2 and CodeLlama show this anomaly on the most imbalanced DiverseVul dataset. It emphasizes the important role of data balance.
Finding 2: The medium-size sequence models excel on short sample datasets, generally outperforming LLMs. On the ReVeal, Draper, and DiverseVul, both of the medium-size sequence models achieve the highest and second-highest F1-scores respectively. There is an evident performance difference between the LLM and medium-sized sequence models. Although they have generally similar precisions, LLMs’ recall rates are significantly lower than the medium-sized sequence models, resulting in lower F1- scores. This gap narrows as the dataset becomes more balanced, with the Devign dataset showing the smallest gap. We conclude that medium-sized sequence models are less affected by a low proportion of positive samples than LLMs. The code pre-trained Transformer encoder enables them to capture vulnerability features more accurately even with limited vulnerability data.
@@ VI-B Analysis on Main Experiments
From the above analysis, it is clear that the proportion of positive samples in the dataset plays a decisive role in the CVD performance of trained models, while the impact of sample length should not be overlooked. To specifically investigate the effects of positive sample ratio and length on fine-tuning LLMs, we have designed experiments in Section VI-C and VI-D respectively.
@@
Figure 5: Metrics on Varying Code Sequence Length.
VI-C Analysis on Datasets with Varying Postive Sample Ratios
As two recent study [4, 5] and our statistics in Table II point out, long-tailed distribution within CVD datasets could pose a challenge for LLMs-based vulnerability detection solutions, and we can also observe this in our main experiments in Table III. Thus, we carry out the re-sampling experiments which creates more balanced datasets. We subsample the Draper and DiverseVul datasets [26, 22] to make the positive samples (i.e. vulnerable) to occupy more percentage across the training dataset, and the ratio is incrementally set to 10%, 20%, 30%, 40% and 50%. Similar to the main experiment, the size of each sampled dataset is controlled at 25,000. We select CodeBERT as the studied medium-size sequence model and the LLama-3.1 as the studied LLM.
Figure 6: Sensitivity Study on LoRA Rank
Figure 7: Sensitivity Study on LoRA Scaling Factor
VI-D Analysis on Analysis on Datasets with Varying Sample Lengths